OpenStudy (anonymous):

63/54 in simplest form

OpenStudy (anonymous):

7/6 bc 9 goes into them both and is the highest number that can.

OpenStudy (amoodarya):

63=9*7 54=9*6 63 7 --=-- 54 6

OpenStudy (anonymous):

no, more like 

OpenStudy (anonymous):

@amoodarya